Circular No. 5883

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SGR 1806–20 B. A. Cooke, Department of Physics and Astronomy, Leicester Univer-

sity, reports: “The x-ray source reported by Y. Tanaka on IAUC 5880 is evident in a much weaker state (0.003 mCrab at 0.22-2.5 keV) in a ROSAT XRT PSPC observation in April. The source position is α = 18^{h}05^{m}41^{s}.56, δ = −20^{o}25′03′′.4 (equinox 1950.0), with an error box of radius 11′′. The box includes the ‘compact nebula’ reported by Kulkarni et al. on IAUC 5879. This further strengthens the connection between SGR 1806–20, the radio source, and the highly variable x-ray source.”

S. Kulkarni, K. Matthews, G. Neugebauer, N. Reid, T. Soifer, and G. Vasisht, California Institute of Technology, communicate: “Observations at the Hale 5-m telescope on Oct. 8 UT reveal a highly reddened star in the vicinity of the peak of the compact nebula that is most likely associated with SGR 1806–20 (cf. IAUC 5879, 5880), the star having the following magnitudes: J = 13.3, H = 10.2, K = 8.85, and [3.4 μm] = 7.2. The position of this object was directly measured with the telescope using the three nearest SAOC stars as references. The star is 2′′.2 west and 1′′.2 north of the radio peak of the compact nebula; the 1σ uncertainty is 1′′. I-band images were obtained with a CCD camera at the Palomar 1.5-m reflector on Oct. 6.11; a star of I = 21 is seen coincident with the radio peak within the astrometric uncertainty (1σ) of 0′′.8. In addition, another star (I = 19) is seen 3′′.5 west and 1′′.3 north of the radio peak; we presume this to be the reddened star discussed above and hence conclude that it is not associated with the compact radio nebula. In view of the recent activity of the Soft Gamma Repeater (cf. IAUC 5875), we urge further monitoring of the fainter star.”

SUPERNOVAE 1977I, 1978L, 1993Y, 1993Z D. D. Balam, University of Victoria, reports his astrometry of SN

1993Y in UGC 2771 and SN 1993Z in NGC 2775 (cf. IAUC 5870), obtained at Climenhaga Observatory (equinox 1950.0; mean errors are 0′′.3 and 0′′.5, respectively): SN 1993Y, α = 3^{h}28^{m}06^{s}.75, δ = +39^{o}34′51′′.9; SN 1993Z, α = 9^{h}07^{m}40^{s}.18, δ = +7^{o}13′56′′.1.

Corrigendum. On IAUC 5682, it was SN 1977I, not SN 1978L, that was located in ESO 412-G6 = MCG –05-03-026.

1993 October 25 Daniel W. E. Green